# Multispecies inhomogeneous $t$-PushTASEP from antisymmetric fusion

### v2

@@ 4.1. Definition
Recall that $\mathbb{V}={\mathsf{V}}^{\otimes L}$ and we have identified ${\mathsf{V}}$ with $V^{1}$ as in (15). Define the transfer matrix $T^{k}(z)=T^{k}(z|x_{1},\ldots,x_{L}):\mathbb{V}\longrightarrow\mathbb{V}$ on the length $L$ periodic lattice by
$T^{k}(z)=\mathrm{Tr}_{V^{k}}\left(S_{0,L}\Bigl{(}\frac{z}{x_{L}}\Bigr{)}\cdots S_{0,1}\Bigl{(}\frac{z}{x_{1}}\Bigr{)}\right)\qquad(0\leq k\leq n+1),$ (35)
where the index 0 denotes the auxiliary space $V^{k}$ over which the trace is taken. The factor $S_{0,r}(z/x_{r})$ is the matrix $S^{k,1}(z/x_{r})$ defined by (29) and (33), which acts on $V^{k}\otimes(\text{$r$'th component of $\mathbb{V}$ from the left})$. Explicitly, one has
$T^{k}(z)|\sigma_{1},\ldots,\sigma_{L}\rangle$ $=\sum_{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\in\{0,\ldots,n\}}T^{k}(z)^{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}}_{\sigma_{1},\ldots,\sigma_{L}}|\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\rangle,$ (36a)
$T^{k}(z)^{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}}_{\sigma_{1},\ldots,\sigma_{L}}$ $=\sum_{{\bf a}_{1},\ldots,{\bf a}_{L}\in\mathscr{B}^{k}}S\Bigl{(}\frac{z}{x_{1}}\Bigr{)}^{{\bf a}_{2},{\bf e}_{\sigma^{\prime}_{1}}}_{{\bf a}_{1},{\bf e}_{\sigma_{1}}}S\Bigl{(}\frac{z}{x_{2}}\Bigr{)}^{{\bf a}_{3},{\bf e}_{\sigma^{\prime}_{2}}}_{{\bf a}_{2},{\bf e}_{\sigma_{2}}}\cdots S\Bigl{(}\frac{z}{x_{L}}\Bigr{)}^{{\bf a}_{1},{\bf e}_{\sigma^{\prime}_{L}}}_{{\bf a}_{L},{\bf e}_{\sigma_{L}}}.$ (36b)
We write the element (36b) as $\langle\boldsymbol{\sigma}^{\prime}|T^{k}(z)|\boldsymbol{\sigma}\rangle$, and depict it as Figure 2.
@@
The parameter $z$ is referred to as the spectral parameter, while $x_{1},\ldots,x_{L}$ represents the inhomogeneity associated with the vertices. Adopting the terminology from the box-ball systems [IKT12], we refer to the ${\bf a}_{1},\ldots,{\bf a}_{L}\in\mathscr{B}^{k}$ as carriers with capacity $k$.
@@ 4.2. Basic properties
From the Yang-Baxter relation (28) with $(k_{1},k_{2},k_{3})=(k,k^{\prime},1)$, one can show the commutativity
$[T^{k}(z|x_{1},\ldots,x_{L}),T^{k^{\prime}}(z^{\prime}|x_{1},\ldots,x_{L})]=0\qquad(0\leq k,k^{\prime}\leq n+1).$ (37)
It is essential to choose the inhomogeneities $x_{1},\ldots,x_{L}$ in the two transfer matrices identically. From the weight conservation property of $S^{k,1}(z)$ and the periodic boundary condition, $T^{k}(z)$ preserves each sector $\mathbb{V}({\bf m})$ in (5).
Let us examine the diagonal elements of $T^{k}(z)$ for general $k\in\{0,\ldots,n+1\}$. When $\boldsymbol{\sigma}^{\prime}=\boldsymbol{\sigma}$, all the arrays ${\bf a}_{j}$ in Figure 2 become identical due to the weight conservation. Thus, by employing (33), we obtain
$\langle\boldsymbol{\sigma}|T^{k}(z)|\boldsymbol{\sigma}\rangle=\sum_{{\bf a}\in\mathscr{B}^{k}}\prod_{j=1}^{L}S\Bigl{(}\frac{z}{x_{j}}\Bigr{)}^{{\bf a},{\bf e}_{\sigma_{j}}}_{{\bf a},{\bf e}_{\sigma_{j}}}=\sum_{{\bf a}\in\mathscr{B}^{k}}\prod_{j=1}^{L}t^{a_{1+\sigma_{j}}+\cdots+a_{n}}\left(1-t^{a_{\sigma_{j}}}\frac{z}{x_{j}}\right).$ (38)
In the special cases of $k=0$ and $n+1$, one has $\mathscr{B}^{0}=\{{\bf 0}:=(0,\ldots,0)\}$ and $\mathscr{B}^{n+1}=\{{\bf 1}:=(1,\ldots,1)\}$ in the multiplicity representation (13). Then, the RHS of (33) becomes $[{\bf a}={\bf i}={\bf 0},j=b](1-z)$ for $k=0$ and $[{\bf a}={\bf i}={\bf 1},j=b]t^{n-j}(1-tz)$ for $k=n+1$. This implies that $T^{0}(z)$ and $T^{n+1}(z)$ are diagonal, with their elements obtained by reducing the sum (38) to the terms ${\bf a}={\bf 0}$ and ${\bf 1}$, respectively. Consequently we have
$T^{0}(z)$ $=\prod_{j=1}^{L}\left(1-\frac{z}{x_{j}}\right)\mathrm{Id},$ (39)
$T^{n+1}(z)$ $=t^{K_{1}+\cdots+K_{n}}\prod_{j=1}^{L}\left(1-\frac{tz}{x_{j}}\right)\mathrm{Id}$ (40)
on $\mathbb{V}({\bf m})$, where $K_{i}$ is defined in (7).
Suppose that we are in the sector $T^{k}(z)\in\mathrm{End}(\mathbb{V}({\bf m}))$ (5), hence $\boldsymbol{\sigma}=(\sigma_{1},\ldots,\sigma_{L})\in\mathcal{S}({\bf m})$ as in (6). From (38), we have
$\begin{split}\langle\boldsymbol{\sigma}|T^{k}(0)|\boldsymbol{\sigma}\rangle&=\sum_{{\bf a}\in\mathscr{B}^{k}}t^{\sum_{j=1}^{L}(a_{1+\sigma_{j}}+\cdots+a_{n})}\\ &=\sum_{{\bf a}\in\mathscr{B}^{k}}t^{m_{0}(a_{1}+\cdots+a_{n})+m_{1}(a_{2}+\cdots+a_{n})+\cdots+m_{n-1}a_{n}}\\ &=\sum_{{\bf a}\in\mathscr{B}^{k}}t^{m_{0}a_{1}+(m_{0}+m_{1})a_{2}+\cdots+(m_{0}+\cdots+m_{n-1})a_{n}}.\end{split}$ (41)
@@
$\langle\boldsymbol{\sigma}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle=-\sum_{j=1}^{L}\frac{1}{x_{j}}\sum_{{\bf a}\in\mathscr{B}^{k}}t^{a_{\sigma_{j}}+m_{0}a_{1}+(m_{0}+m_{1})a_{2}+\cdots+(m_{0}+\cdots+m_{n-1})a_{n}}.$ (42)
@@
$\langle\boldsymbol{\sigma}|T^{k}(0)|\boldsymbol{\sigma}\rangle$ $=e_{k}(t^{K_{0}},\ldots,t^{K_{n}}),\qquad$ (43)
$\langle\boldsymbol{\sigma}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle$ $=-\sum_{j=1}^{L}\frac{1}{x_{j}}e_{k}(u^{(\sigma_{j})}_{0},\ldots,u^{(\sigma_{j})}_{n}),\qquad u^{(\sigma)}_{i}=t^{\delta_{i,\sigma}+K_{i}}.$ (44)
where $K_{i}$ is defined in (7). In particular $u_{0}^{(\sigma)}=t^{\delta_{0,\sigma}}$. The functions $e_{0},\ldots,e_{n+1}$ are elementary symmetric polynomials in $(n+1)$ variables defined by
$e_{k}(w_{0},\ldots,w_{n})=\sum_{{\bf a}\in\mathscr{B}^{k}}w_{0}^{a_{0}}\cdots w_{n}^{a_{n}},$ (45)
@@
$(1+\zeta w_{0})\cdots(1+\zeta w_{n})=\sum_{k=0}^{n+1}\zeta^{k}e_{k}(w_{0},\ldots,w_{n}).$ (46)
@@
$\sum_{k=0}^{n+1}(-1)^{k}e_{k}(u^{(\sigma)}_{0},\ldots,u^{(\sigma)}_{n})$ $=(1-u^{(\sigma)}_{0})\cdots(1-u^{(\sigma)}_{n})=\delta_{\sigma,0}D_{\bf m}\qquad(0\leq\sigma\leq n),$ (47)
where $D_{\bf m}$ is defined in (8).
From the definition (36a)–(36b) and (33), the transfer matrix $T^{k}(z)$ is diagonal at $z=0$. Therefore, (43) implies
$T^{k}(0)=e_{k}(t^{K_{0}},\ldots,t^{K_{n}})\mathrm{Id}\;\;\text{on }\mathbb{V}({\bf m}).$ (48)
Example 2.
Following Example 1, we set $n=2$ and $L=4$. We denote the coefficient of the diagonal term generated by $T^{k}(z)$ by $\mathcal{D}_{k}(z)$.
$\begin{split}T^{0}(z)|0121\rangle&=\mathcal{D}_{0}(z)|0121\rangle,\\ T^{1}(z)|0121\rangle&=\frac{(1-t)^{4}z^{2}}{x_{2}x_{3}}|1012\rangle-\frac{(1-t)^{3}z(z-x_{2})}{x_{2}x_{3}}|1102\rangle+\frac{(1-t)^{2}tz(z-x_{1})(tz-x_{2})}{x_{1}x_{2}x_{3}}|0112\rangle\\ &+\frac{(1-t)^{2}z(z-x_{2})(z-x_{3})}{x_{2}x_{3}x_{4}}|1120\rangle+\frac{(1-t)^{2}t^{2}z(z-x_{1})(z-x_{4})}{x_{1}x_{3}x_{4}}|0211\rangle\\ &-\frac{(1-t)^{3}tz^{2}(z-x_{4})}{x_{2}x_{3}x_{4}}|2011\rangle+\frac{(1-t)^{2}tz(z-x_{2})(z-x_{4})}{x_{2}x_{3}x_{4}}|2101\rangle\\ &+\frac{(1-t)^{2}z(z-x_{3})(tz-x_{4})}{x_{2}x_{3}x_{4}}|1021\rangle+\mathcal{D}_{1}(z)|0121\rangle,\\ T^{2}(z)|0121\rangle&=\frac{(1-t)^{2}tz(tz-x_{1})(tz-x_{2})}{x_{1}x_{2}x_{3}}|0112\rangle+\frac{(1-t)^{4}t^{2}z^{2}}{x_{3}x_{4}}|1210\rangle\\ &+\frac{(1-t)^{3}t^{2}z^{2}(tz-x_{2})}{x_{2}x_{3}x_{4}}|2110\rangle+\frac{(1-t)^{2}t^{3}z(z-x_{2})(tz-x_{3})}{x_{2}x_{3}x_{4}}|1120\rangle\\ &+\frac{(1-t)^{2}t^{2}z(tz-x_{1})(z-x_{4})}{x_{1}x_{3}x_{4}}|0211\rangle+\frac{(1-t)^{3}t^{2}z(tz-x_{4})}{x_{3}x_{4}}|1201\rangle\\ &+\frac{(1-t)^{2}t^{2}z(tz-x_{2})(tz-x_{4})}{x_{2}x_{3}x_{4}}|2101\rangle+\frac{(1-t)^{2}t^{3}z(tz-x_{3})(tz-x_{4})}{x_{2}x_{3}x_{4}}|1021\rangle+\mathcal{D}_{2}(z)|0121\rangle,\\ T^{3}(z)|0121\rangle&=\mathcal{D}_{3}(z)|0121\rangle.\end{split}$ (49)
The functions $\mathcal{D}_{0}(z)$ and $\mathcal{D}_{3}(z)$ are explicitly given by (39) and (40) with $n=2$, respectively. They lead to
$\begin{split}\dot{T}^{0}(0)|0121\rangle&=\dot{\mathcal{D}}_{0}(0)|0121\rangle,\\ \dot{T}^{1}(0)|0121\rangle&=\frac{(1-t)^{2}}{x_{2}}|1021\rangle+\frac{(1-t)^{2}t}{x_{3}}|0112\rangle+\frac{(1-t)^{2}t^{2}}{x_{3}}|0211\rangle\\ &+\frac{(1-t)^{3}}{x_{3}}|1102\rangle+\frac{(1-t)^{2}t}{x_{3}}|2101\rangle+\frac{(1-t)^{2}}{x_{4}}|1120\rangle+\dot{\mathcal{D}}_{1}(0)|0121\rangle,\\ \dot{T}^{2}(0)|0121\rangle&=\frac{(1-t)^{2}t^{3}}{x_{2}}|1021\rangle+\frac{(1-t)^{2}t}{x_{3}}|0112\rangle+\frac{(1-t)^{2}t^{2}}{x_{3}}|0211\rangle\\ &-\frac{(1-t)^{3}t^{2}}{x_{3}}|1201\rangle+\frac{(1-t)^{2}t^{2}}{x_{3}}|2101\rangle+\frac{(1-t)^{2}t^{3}}{x_{4}}|1120\rangle+\dot{\mathcal{D}}_{2}(0)|0121\rangle,\\ \dot{T}^{3}(0)|0121\rangle&=\dot{\mathcal{D}}_{3}(0)|0121\rangle,\end{split}$ (50)
where $\dot{\mathcal{D}}_{k}(0)=\left.\frac{d\mathcal{D}_{k}(z)}{dz}\right|_{z=0}$ is available from (44).

### v3

@@ 4.1. Definition
Recall that $\mathbb{V}={\mathsf{V}}^{\otimes L}$ and we have identified $\mathsf{V}$} with $V^{1}$ as in (3.3). Define the transfer matrix $T^{k}(z)=T^{k}(z|x_{1},\ldots,x_{L}):\mathbb{V}\longrightarrow\mathbb{V}$ on the length $L$ periodic lattice by
$T^{k}(z)=\mathrm{Tr}_{V^{k}}\left(S_{0,L}\Bigl(\frac{z}{x_{L}}\Bigr)\cdots S_{0,1}\Bigl(\frac{z}{x_{1}}\Bigr)\right)\qquad(0\leq k\leq n+1),$ (4.1)
where the index 0 denotes the auxiliary space $V^{k}$ over which the trace is taken. The factor $S_{0,r}(z/x_{r})$ is the matrix $S^{k,1}(z/x_{r})$ defined by (3.5) and (3.12), which acts on $V^{k}\otimes(\text{$r$'th component of $\mathbb{V}$ from the left})$. Explicitly, one has
$T^{k}(z)|\sigma_{1},\ldots,\sigma_{L}\rangle$ $=\sum_{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\in\{0,\ldots,n\}}T^{k}(z)^{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}}_{\sigma_{1},\ldots,\sigma_{L}}|\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}\rangle,$ (4.2a)
$T^{k}(z)^{\sigma^{\prime}_{1},\ldots,\sigma^{\prime}_{L}}_{\sigma_{1},\ldots,\sigma_{L}}$ $=\sum_{{\bf a}_{1},\ldots,{\bf a}_{L}\in\mathscr{B}^{k}}S\Bigl(\frac{z}{x_{1}}\Bigr)^{{\bf a}_{2},{\bf e}_{\sigma^{\prime}_{1}}}_{{\bf a}_{1},{\bf e}_{\sigma_{1}}}S\Bigl(\frac{z}{x_{2}}\Bigr)^{{\bf a}_{3},{\bf e}_{\sigma^{\prime}_{2}}}_{{\bf a}_{2},{\bf e}_{\sigma_{2}}}\cdots S\Bigl(\frac{z}{x_{L}}\Bigr)^{{\bf a}_{1},{\bf e}_{\sigma^{\prime}_{L}}}_{{\bf a}_{L},{\bf e}_{\sigma_{L}}}.$ (4.2b)
We write the element (4.2b) as $\langle\boldsymbol{\sigma}^{\prime}|T^{k}(z)|\boldsymbol{\sigma}\rangle$, and depict it as Figure 5.
@@
The parameter $z$ is referred to as the spectral parameter, while $x_{1},\ldots,x_{L}$ represent the inhomogeneity associated with the vertices. Adopting the terminology from the box-ball systems [IKT12], we refer to the ${\bf a}_{1},\ldots,{\bf a}_{L}\in\mathscr{B}^{k}$ as carriers with capacity $k$.
@@ 4.2. Basic properties
From the Yang-Baxter relation (3.13) with $(k_{1},k_{2},k_{3})=(k,k^{\prime},1)$, one can show the commutativity
$[T^{k}(z|x_{1},\ldots,x_{L}),T^{k^{\prime}}(z^{\prime}|x_{1},\ldots,x_{L})]=0\qquad(0\leq k,k^{\prime}\leq n+1).$ (4.3)
It is essential to choose the inhomogeneities $x_{1},\ldots,x_{L}$ in the two transfer matrices identically. From the weight conservation property of $S^{k,1}(z)$ and the periodic boundary condition, $T^{k}(z)$ preserves each sector $\mathbb{V}({\bf m})$ in (2.1).
Let us examine the diagonal elements of $T^{k}(z)$ for general $k\in\{0,\ldots,n+1\}$. When $\boldsymbol{\sigma}^{\prime}=\boldsymbol{\sigma}$, all the arrays ${\bf a}_{j}$ in Figure 5 become identical due to the weight conservation. Thus, by employing (3.12), we obtain
$\langle\boldsymbol{\sigma}|T^{k}(z)|\boldsymbol{\sigma}\rangle=\sum_{{\bf a}\in\mathscr{B}^{k}}\prod_{j=1}^{L}S\Bigl(\frac{z}{x_{j}}\Bigr)^{{\bf a},{\bf e}_{\sigma_{j}}}_{{\bf a},{\bf e}_{\sigma_{j}}}=\sum_{{\bf a}\in\mathscr{B}^{k}}\prod_{j=1}^{L}t^{a_{1+\sigma_{j}}+\cdots+a_{n}}\left(1-t^{a_{\sigma_{j}}}\frac{z}{x_{j}}\right).$ (4.4)
In the special cases of $k=0$ and $n+1$, one has $\mathscr{B}^{0}=\{{\bf 0}:=(0,\ldots,0)\}$ and $\mathscr{B}^{n+1}=\{{\bf 1}:=(1,\ldots,1)\}$ in the multiplicity representation (3.1). Then, the RHS of (3.12) becomes ${\bf a}={\bf i}={\bf 0},j=b](1-z)$ for $k=0$ and ${\bf a}={\bf i}={\bf 1},j=b]t^{n-j}(1-tz)$ for $k=n+1$. This implies that $T^{0}(z)$ and $T^{n+1}(z)$ are diagonal, with their elements obtained by reducing the sum (4.4) to the terms ${\bf a}={\bf 0}$ and $\bf 1$}, respectively. Consequently we have
$T^{0}(z)$ $=\prod_{j=1}^{L}\left(1-\frac{z}{x_{j}}\right)\mathrm{Id},$ (4.5)
$T^{n+1}(z)$ $=t^{K_{1}+\cdots+K_{n}}\prod_{j=1}^{L}\left(1-\frac{tz}{x_{j}}\right)\mathrm{Id}$ (4.6)
on $\mathbb{V}({\bf m})$, where $K_{i}$ is defined in (2.3).
Suppose that we are in the sector $T^{k}(z)\in\mathrm{End}(\mathbb{V}({\bf m}))$ (2.1), hence $\boldsymbol{\sigma}=(\sigma_{1},\ldots,\sigma_{L})\in\mathcal{S}({\bf m})$ as in (2.2). From (4.4), we have
$\begin{split}\langle\boldsymbol{\sigma}|T^{k}(0)|\boldsymbol{\sigma}\rangle&=\sum_{{\bf a}\in\mathscr{B}^{k}}t^{\sum_{j=1}^{L}(a_{1+\sigma_{j}}+\cdots+a_{n})}\\ &=\sum_{{\bf a}\in\mathscr{B}^{k}}t^{m_{0}(a_{1}+\cdots+a_{n})+m_{1}(a_{2}+\cdots+a_{n})+\cdots+m_{n-1}a_{n}}\\ &=\sum_{{\bf a}\in\mathscr{B}^{k}}t^{m_{0}a_{1}+(m_{0}+m_{1})a_{2}+\cdots+(m_{0}+\cdots+m_{n-1})a_{n}}.\end{split}$ (4.7)
@@
$\langle\boldsymbol{\sigma}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle=-\sum_{j=1}^{L}\frac{1}{x_{j}}\sum_{{\bf a}\in\mathscr{B}^{k}}t^{a_{\sigma_{j}}+m_{0}a_{1}+(m_{0}+m_{1})a_{2}+\cdots+(m_{0}+\cdots+m_{n-1})a_{n}}.$ (4.8)
@@
$\langle\boldsymbol{\sigma}|T^{k}(0)|\boldsymbol{\sigma}\rangle$ $=e_{k}(t^{K_{0}},\ldots,t^{K_{n}}),\qquad$ (4.9)
$\langle\boldsymbol{\sigma}|\dot{T}^{k}(0)|\boldsymbol{\sigma}\rangle$ $=-\sum_{j=1}^{L}\frac{1}{x_{j}}e_{k}(u^{(\sigma_{j})}_{0},\ldots,u^{(\sigma_{j})}_{n}),\qquad u^{(\sigma)}_{i}=t^{\delta_{i,\sigma}+K_{i}}.$ (4.10)
where $K_{i}$ is defined in (2.3). In particular $u_{0}^{(\sigma)}=t^{\delta_{0,\sigma}}$. The functions $e_{0},\ldots,e_{n+1}$ are elementary symmetric polynomials in $(n+1)$ variables defined by
$e_{k}(w_{0},\ldots,w_{n})=\sum_{{\bf a}\in\mathscr{B}^{k}}w_{0}^{a_{0}}\cdots w_{n}^{a_{n}},$ (4.11)
@@
$(1+\zeta w_{0})\cdots(1+\zeta w_{n})=\sum_{k=0}^{n+1}\zeta^{k}e_{k}(w_{0},\ldots,w_{n}).$ (4.12)
@@
$\sum_{k=0}^{n+1}(-1)^{k}e_{k}(u^{(\sigma)}_{0},\ldots,u^{(\sigma)}_{n})$ $=(1-u^{(\sigma)}_{0})\cdots(1-u^{(\sigma)}_{n})=\delta_{\sigma,0}D_{\bf m}\qquad(0\leq\sigma\leq n),$ (4.13)
where $D_{\bf m}$ is defined in (2.4).
From the definition (4.2a)–(4.2b) and (3.12), the transfer matrix $T^{k}(z)$ is diagonal at $z=0$. Therefore, (4.9) implies
$T^{k}(0)=e_{k}(t^{K_{0}},\ldots,t^{K_{n}})\mathrm{Id}\;\;\text{on }\mathbb{V}({\bf m}).$ (4.14)
Example 4.1.
Following Example 2.1, we set $n=2$ and $L=4$. We denote the coefficient of the diagonal term generated by $T^{k}(z)$ by $\mathcal{D}_{k}(z)$.
$\begin{split}T^{0}(z)|0121\rangle&=\mathcal{D}_{0}(z)|0121\rangle,\\ T^{1}(z)|0121\rangle&=\frac{(1-t)^{4}z^{2}}{x_{2}x_{3}}|1012\rangle-\frac{(1-t)^{3}z(z-x_{2})}{x_{2}x_{3}}|1102\rangle\\ &+\frac{(1-t)^{2}tz(z-x_{1})(tz-x_{2})}{x_{1}x_{2}x_{3}}|0112\rangle+\frac{(1-t)^{2}z(z-x_{2})(z-x_{3})}{x_{2}x_{3}x_{4}}|1120\rangle\\ &+\frac{(1-t)^{2}t^{2}z(z-x_{1})(z-x_{4})}{x_{1}x_{3}x_{4}}|0211\rangle-\frac{(1-t)^{3}tz^{2}(z-x_{4})}{x_{2}x_{3}x_{4}}|2011\rangle\\ &+\frac{(1-t)^{2}tz(z-x_{2})(z-x_{4})}{x_{2}x_{3}x_{4}}|2101\rangle+\frac{(1-t)^{2}z(z-x_{3})(tz-x_{4})}{x_{2}x_{3}x_{4}}|1021\rangle\\ &+\mathcal{D}_{1}(z)|0121\rangle,\\ T^{2}(z)|0121\rangle&=\frac{(1-t)^{2}tz(tz-x_{1})(tz-x_{2})}{x_{1}x_{2}x_{3}}|0112\rangle+\frac{(1-t)^{4}t^{2}z^{2}}{x_{3}x_{4}}|1210\rangle\\ &+\frac{(1-t)^{3}t^{2}z^{2}(tz-x_{2})}{x_{2}x_{3}x_{4}}|2110\rangle+\frac{(1-t)^{2}t^{3}z(z-x_{2})(tz-x_{3})}{x_{2}x_{3}x_{4}}|1120\rangle\\ &+\frac{(1-t)^{2}t^{2}z(tz-x_{1})(z-x_{4})}{x_{1}x_{3}x_{4}}|0211\rangle+\frac{(1-t)^{3}t^{2}z(tz-x_{4})}{x_{3}x_{4}}|1201\rangle\\ &+\frac{(1-t)^{2}t^{2}z(tz-x_{2})(tz-x_{4})}{x_{2}x_{3}x_{4}}|2101\rangle+\frac{(1-t)^{2}t^{3}z(tz-x_{3})(tz-x_{4})}{x_{2}x_{3}x_{4}}|1021\rangle\\ &+\mathcal{D}_{2}(z)|0121\rangle,\\ T^{3}(z)|0121\rangle&=\mathcal{D}_{3}(z)|0121\rangle.\end{split}$ (4.15)
The functions $\mathcal{D}_{0}(z)$ and $\mathcal{D}_{3}(z)$ are explicitly given by (4.5) and (4.6) with $n=2$, respectively. They lead to
$\begin{split}\dot{T}^{0}(0)|0121\rangle&=\dot{\mathcal{D}}_{0}(0)|0121\rangle,\\ \dot{T}^{1}(0)|0121\rangle&=\frac{(1-t)^{2}}{x_{2}}|1021\rangle+\frac{(1-t)^{2}t}{x_{3}}|0112\rangle+\frac{(1-t)^{2}t^{2}}{x_{3}}|0211\rangle\\ &+\frac{(1-t)^{3}}{x_{3}}|1102\rangle+\frac{(1-t)^{2}t}{x_{3}}|2101\rangle+\frac{(1-t)^{2}}{x_{4}}|1120\rangle+\dot{\mathcal{D}}_{1}(0)|0121\rangle,\\ \dot{T}^{2}(0)|0121\rangle&=\frac{(1-t)^{2}t^{3}}{x_{2}}|1021\rangle+\frac{(1-t)^{2}t}{x_{3}}|0112\rangle+\frac{(1-t)^{2}t^{2}}{x_{3}}|0211\rangle\\ &-\frac{(1-t)^{3}t^{2}}{x_{3}}|1201\rangle+\frac{(1-t)^{2}t^{2}}{x_{3}}|2101\rangle+\frac{(1-t)^{2}t^{3}}{x_{4}}|1120\rangle+\dot{\mathcal{D}}_{2}(0)|0121\rangle,\\ \dot{T}^{3}(0)|0121\rangle&=\dot{\mathcal{D}}_{3}(0)|0121\rangle,\end{split}$ (4.16)
where $\dot{\mathcal{D}}_{k}(0)=\left.\frac{d\mathcal{D}_{k}(z)}{dz}\right|_{z=0}$ is available from (4.10).